APPLICATION OF THE MAGNETIC FIELD EFFECTS IN THE GASEOUS FLUORESCENCE AND PHO-TOCHEMISTRY TO INVESTIGATION OF THE PHOTOLYSIS MECHANISMS

\underline{VLADIMIR I. MAKAROV}, Institute of Chemical Kinetics and Combustion SB RAS, 630090, Novosibirsk, Russia; SERGEI A. KOCHUBEI, VALERII N. ISCHENKO, Institute of Semiconductor Physics, SB RAS, 630090, Novosibirsk, Russia.

Since the earlier work of Solarz, Batler and Levy in 1973 on NO_{2} and Matsuzaki and Nagakura in 1974 on CS_{2}, external magnetic field effects on gaseous fluorescence have been found for a number of molecules such as SO_{2}, carbonyls, diazines and triazine. In the present study, we have investigated the magnetic field effects in the NO_{2}, SO_{2}, and (COF)2 (000band fluorescence of the Ã^{1}Au←X̃^{1}Ag transition) fluorescence and in the NO_{2} and CS_{2} photolysis under the time resolved experiments. It was found that in the presence of a magnetic field, the fast component induced by a magnetic field appears in the NO_{2} and SO_{2} fluorescence. Lifetime and amplitude of this component are dependent of a magnetic field (B). The observed data were explained by DM, in a frames of which, we assume that the ⟨^{2}B_{2} |H_{z} |^{2}A_{1}⟩ (= VB) and ⟨^{2}B_{1} |H_{z} |^{2}A_{1}⟩ (=VB) coupling are significant for NO_{2}, and ( ⟨^{1}A_{2} |H_{z} |^{1}A_{1}⟩+( ⟨^{1}B_{1} |H_{z} |^{1}A_{1}⟩ (= VB = bB) coupling is significant for SO_{2}. Here, Hz is the Zeeman operator. It has also been found that the magnetic field reduces the yield of the NO_{3} radical and S_{2} molecules under the NO_{2} and CS_{2} photolysis. From analysis of these data it has been defined the effective reactivity of the different excited states of the studied systems. It was found that the fast component of the 00-level fluorescence decay of the Ã^{1}Au←X̃^{1}Ag transition appears in the presence of a magnetic field. Lifetime of this component is not dependent of B, while it amplitude increases with increasing of B. Lifetime of the slow component increases in the presence of a magnetic field at P((COF)2) = 0.3 mTorr, while it decreases under other pressures. Obtained data were explained by IM, ie, magnetic field induces coupling of the hyperfine and fine sublevels of the neighbouring triplet state, which are coupled by the intramolecular interactions with the fluorescent singlet levels. Since in this case, a magnetic field influences on the efficiency of the S-T coversion, using of this method we can study the role of such energy transformation in the (COF)2 molecule photolysis.